OpenStudy (anonymous):

A landscaper earns $30 for each lawn her company mows, but she pays $210 per day in salary to her employees. If her company made more than $150 profit from mowing lawns in a 7-day week, what are the possible numbers of lawns the company could have mowed? Check all that apply. 0 5 7 12 37 54 61 80

OpenStudy (anonymous):

@aloud help he cuhh

OpenStudy (anonymous):

30x-210y=150 Let's solve for x. 30x−210y=150 Step 1: Add 210y to both sides. 30x−210y+210y=150+210y 30x=210y+150 Step 2: Divide both sides by 30. 30x30=210y+15030 x=7y+5: i think this is it but i might of made a few err you might wont to get it checked

OpenStudy (anonymous):

hey @mathmate can you help

OpenStudy (mathmate):

@aloud In solving word problems, whenever we use a variable not defined in the question, we need to define them, including units, before we set up equations. Let x=number of lawns mowed in 7 days, then 30x-7(210)\(\ge\)150 30x\(\ge\)1470+150 30x\(\ge\)1620 x\(\ge\)54 So in fact, your answer is correct by setting y=7 days.